Suppose that x liters of 30% acid solution are mixed with y liters of a 43% solution to obtain 100 L of a 40% solution. One equation is a system for solving this problem is x+y = 100. Which of the following is the other equation?
a. 0.43x+0.30y = 0.40*100
b. 30x+43y = 40
c. 0.30x+0.43y = 0.40*100
d. 43x+30y = 0.40*100

Answer by mananth(16946)

x+y = 100.................1
...
c. 0.30x+0.43y = 0.40*100
...
m.ananth@hotmail.ca